Vortex Company operates a retail store with two departments. Information about those departments follows.

Department ADepartment B

Sales$838,500 $451,500

Cost of goods sold 440,000  293,475

Direct expenses      

Salaries 115,000  83,000

Insurance 19,000  10,700

Utilities 22,000  28,500

Depreciation 15,500  13,500

Maintenance 6,200  5,300


The company also incurred the following indirect costs.

  

Salaries$31,000

Insurance7,200

Depreciation15,400

Office expenses41,000


Indirect costs are allocated as follows: salaries on the basis of sales; insurance and depreciation on the basis of square footage; and office expenses on the basis of number of employees. Additional information about the departments follows.

DepartmentSquare footageNumber of employees

A 30,800  84

B 13,200  36


Required:
1. Determine the departmental contribution to overhead and the departmental net income for department A and Department B. (Do not round intermediate calculations. Round your final answers to nearest whole dollar.)




2. Should Department B be eliminated?

Yes

No
Answer #1
Department A Department B
Sales $838500 $451500
Less: Cost of goods sold (440000) (293475)
Gross profit 398500 158025
Less: Direct expenses
Salaries 115000 83000
Insurance 19000 10700
Utilities 22000 28500
Depreciation 15500 13500
Maintenance 6200 5300
Total direct expenses 177700 141000
Departmental contribution overhead 220800 17025
Less: Indirect expenses
Salaries 20150 10850
Insurance 5040 2160
Depreciation 10780 4620
Office expenses 28700 12300
Total indirect expenses 64670 29930
Operating income (loss) $156130 $(12905)

Allocation of indirect costs

Total sales= $838500+451500= $1290000

Salaries

Department A= $31000*838500/1290000= $20150

Department B= $31000*451500/1290000= $10850

Insurance

Total Square footage= 30800+13200= 44000

Department A= $7200*30800/44000= $5040

Department B= $7200*13200/44000= $10850

Depreciation

Department A= $15400*30800/44000= $10780

Department B= $15400*13200/44000= $4620

Office expenses

Total number of employees= 84+36= 120

Department A= $41000*84/120= $28700

Department B= $41000*36/120= $12300

2) No, Department B should not be eliminated because Department B is contributing $17025 toward indirect cost and if Department B is eliminated all the indirect cost will be contributed toward Department A by which operating income of Department A will decrease. So, Department B should not be eliminated.